# Multilingual large language models leak human stereotypes across language boundaries

### v2

@@ 3.2. Qualitative Results
Moving forward, we delve into the specific stereotypical associations that leak from one language to another, considering the potential implications of such strengthened associations. We focus on the GPT-3.5 model, in which we observe the most leakage from human stereotypes. Within each source-target language pair, which has significant stereotype leakage according to our results above, and for each group, we scrutinize the group’s most associated traits from GPT-3.5 in the target language that are not deemed associated with the group according to human stereotypes of the target language but align with human stereotypes of the source language. Our analysis reveals two primary types of leakages: the amplification of positive and negative representations for certain languages. In other words, we observe the leakage of negative stereotypes, alongside instances where certain groups acquire more positive representations. Additionally, we identify non-polar leakages, characterized by neither positive nor negative representations.
@@ 3.2.1. Positive Leakage
According to human annotation, “Asian people” are more positively perceived in English language than in Russian. We observe the strengthening of such traits in GPT-3.5 Russian language as wealthy, likable, and high status. Moreover, “housewives” become more warm in English following leakages from Russian and Hindi. “Black people” are more powerful, modern, confident, and wealthy in the English language following leakage from Hindi. Another example of the leakage of positive perceptions is for “gay men” and “lesbians” from English to other languages. Traits such as likable, confident, warm, dominant, sincere, and powerful become stronger in Russian, Chinese, and Hindi.
@@ 3.2.2. Negative Leakage
On the other hand, there are negative stereotypes that leak across languages. From “feminists”, we observe a leakage from English to Chinese and Hindi and from Russian to Chinese of such stereotypical associations as egoistic, threatening, repellent, and cold, while, for instance, in the human data in Hindi this group is perceived as warm.
Another example is “immigrants”. From Russian and English languages, traits such as threatening, repellent, dishonest, egoistic, and unconfident leak to Chinese and Hindi. Based on human data, we found that people surveyed in Chinese view this group quite favorably since the majority of immigrants to China are highly qualified professionals (Pieke, 2012). In Russia, immigrants are mostly coming from poorer neighboring countries and are negatively stereotyped in society, while in the U.S., immigrants are diverse and could be both marginalized or privileged.
@@ 3.2.3. Non-polar Leakage
There are also non-polar leakages, which are neither positive nor negative. From Hindi to English and Russian, we see the strengthening of religious for various groups such as “women”, “disabled people”, “Black people”, and “Asian people”. It has been shown that there are more than believers of the total population in India as of (Sahgal et al., 2021).
@@ 3.2.4. Non-shared Groups Leakage
In the case of non-shared groups, we expected uni-directional transferring of the groups’ perceptions from the language of origin to other languages. Our findings confirm this hypothesis. For example, the group “VDV soldiers” is a widely known military unit in Russia. There are strong stereotypes in Russian society about this group, but the group is mostly unknown to Americans. Out of the survey English survey respondents who passed the quality tests, no one chose this group as a familiar one. This group’s representation leaks from Russian to English, strengthening traits such as confident, traditional, competitive, and threatening. Another example is “Hui people”, a group widely unknown to Russian and Hindi society: out of respondents for both surveys, no one chose this group as the familiar one. This social group is a minority in China and is composed of Chinese-speaking followers of Islam. Originally, “Hui people” are marginalized in China and viewed as more traditional, religious, and conservative (Hillman, 2004; Hong, 2005). Accordingly, we observed the leakage of such traits as irrational, traditional, threatening, repellent, religious, and egoistic. All groups specific to the Hindi language — “Gujarati, Brahmin”, and “Shudra people” — have certain traits leaking to the English and Russian languages. For example, high caste groups (“Gujarati” and “Brahmin people”) strengthen such positive traits as wealthy, likable, sincere, powerful, high status, competitive, and confident. In addition, “Brahmin people” become more associated in GPT-3.5 with traits poor, low status, powerless, traditional, religious, and dominated. This leakage corresponds to the perception of these groups in Indian society and by our survey respondents (Witzel, 1993; Milner, 1993).
3.2.5. Discussion
@@
However, positive stereotypes, such as the notion that “Asian people” are wealthy or “housewives” are warm, can also lead to unrealistic expectations and pressures. For instance, not all Asian people are wealthy, and assuming so can ignore the diverse economic realities faced by individuals within this broad demographic. Similarly, the stereotype that housewives are inherently warm can enforce restrictive roles based on gender.
@@
The leakage of stereotypes is particularly troubling for certain applications, including education and creative content generation. These fields heavily influence public perception and personal development, making the integrity of the content they deliver crucial. Systems built for these applications with MLLMs must be particularly cautious of the stereotype leakage effect. It is essential for developers to implement strategies that actively mitigate the harmful leakage effects.
4. Conclusion & Limitations
@@
Multilingual large language models have the potential to spread stereotypes beyond the societal context they emerge from, whether by generating new stereotypes, amplifying existing ones, or reinforcing prevailing social perceptions from dominant cultures. In our work, we demonstrate that this concern is indeed valid. To do so, we establish a framework for measuring the leakage of stereotypical associations in multilingual large language models across languages. Overall, we find that the stereotype leakage occurs bidirectionally meaning that when one language transmits stereotypes to others, it likely receives some stereotypes from other languages as well. We also observe the most stereotype leakage effect within the GPT-3.5 model. Within the GPT-3.5 model, we observe the strengthening of positive, negative, and non-polar associations in the model. In addition, our study underscores the role of “native” languages in framing social groups unknown to other linguistic communities. Such leakage of stereotypes amplifies the complexity of societal perceptions by introducing a complex interconnected bias from different languages and cultures. In the context of shared groups, stereotype leakage may manifest as the manifestation of stereotypes that were not previously present within the cultural setting of a particular group. In the case of non-shared groups, stereotype leakage can extend the reach of existing stereotypes from the source culture to other cultural contexts.
@@
To our knowledge, we are the first to introduce the concept of stereotype leakage across languages in multilingual LLMs. We propose a framework for quantifying this leakage in multilingual models, which can be easily applied to unstudied social groups. We show that multilingual large language models could facilitate the transmission of biases across different cultures and languages. We demonstrate the existence of stereotype leakage within MLLMs, which are trained on diverse linguistic datasets. As multilingual models begin to play an increasingly influential role in AI applications and across societies, understanding their potential vulnerabilities and the level of bias propagation across linguistic boundaries becomes important. As a result, we lay the groundwork for advancing both the theoretical comprehension of multilingual models and the practical implementation for bias mitigation in AI systems.
Limitations
Our work has several limitations. First, we are limited in our ability to run a causal analysis because none of the studied languages can be easily removed from the training data to see their genuine impact on stereotypical associations in other languages. Retraining GPT-3.5, for instance, is not a feasible option. Thus, we use BERT monolingual model as a proxy for each language. In addition, stereotype traits were selected based on the ABC model, which was developed and tested using U.S. and German stereotypes. Though we translated our surveys into all four languages, the stereotype traits may better reflect Anglocentric stereotypes (Talat et al., 2022) than others. Furthermore, the human stereotypes we collected may already reflect the influence of social stereotype transmission. For instance, in our study, we surveyed crowd workers about their consumption of U.S. social media. We found that, on average, of respondents from Russia, China, and India engage with U.S. social platforms. Such American cultural dominance could potentially affect the human stereotypes collected in these three languages. Lastly, while we indirectly consider culture through survey results on associations, we do not measure or account for culture in a comprehensive manner. Our English language survey results only apply to the U.S., Russian to Russia, Chinese to China, and Hindi to India. Lastly,
@@ Acknowledgements.
- Brown et al. (2020) Tom B. Brown, Benjamin Mann, Nick Ryder, Melanie Subbiah, Jared Kaplan, Prafulla Dhariwal, Arvind Neelakantan, Pranav Shyam, Girish Sastry, Amanda Askell, Sandhini Agarwal, Ariel Herbert-Voss, Gretchen Krueger, Tom Henighan, Rewon Child, Aditya Ramesh, Daniel M. Ziegler, Jeffrey Wu, Clemens Winter, Christopher Hesse, Mark Chen, Eric Sigler, Mateusz Litwin, Scott Gray, Benjamin Chess, Jack Clark, Christopher Berner, Sam McCandlish, Alec Radford, Ilya Sutskever, and Dario Amodei. 2020. Language Models are Few-Shot Learners. CoRR abs/2005.14165 (2020). arXiv:2005.14165 https://arxiv.org/abs/2005.14165
- Cabello Piqueras and Søgaard (2022) Laura Cabello Piqueras and Anders Søgaard. 2022. Are Pretrained Multilingual Models Equally Fair across Languages?. In Proceedings of the 29th International Conference on Computational Linguistics. International Committee on Computational Linguistics, Gyeongju, Republic of Korea, 3597–3605. https://aclanthology.org/2022.coling-1.318
@@
- Choudhury and Deshpande (2021) Monojit Choudhury and Amit Deshpande. 2021. How Linguistically Fair Are Multilingual Pre-Trained Language Models? Proceedings of the AAAI Conference on Artificial Intelligence 35, 14 (May 2021), 12710–12718. https://doi.org/10.1609/aaai.v35i14.17505
@@
- Goldstein et al. (2023) Josh A. Goldstein, Girish Sastry, Micah Musser, Renee DiResta, Matthew Gentzel, and Katerina Sedova. 2023. Generative Language Models and Automated Influence Operations: Emerging Threats and Potential Mitigations. arXiv:2301.04246 [cs.CY]
- González et al. (2020) Ana Valeria González, Maria Barrett, Rasmus Hvingelby, Kellie Webster, and Anders Søgaard. 2020. Type B Reflexivization as an Unambiguous Testbed for Multilingual Multi-Task Gender Bias. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 2637–2648. https://doi.org/10.18653/v1/2020.emnlp-main.209
@@
- K et al. (2020) Karthikeyan K, Zihan Wang, Stephen Mayhew, and Dan Roth. 2020. Cross-Lingual Ability of Multilingual BERT: An Empirical Study. In International Conference on Learning Representations. https://openreview.net/forum?id=HJeT3yrtDr
@@
- Koch et al. (2016) Alex Koch, Roland Imhoff, Ron Dotsch, Christian Unkelbach, and Hans Alves. 2016. The ABC of Stereotypes About Groups: Agency/Socioeconomic Success, Conservative-Progressive Beliefs, and Communion. Journal of personality and social psychology 110 (05 2016), 675–709. https://doi.org/10.1037/pspa0000046
@@
- Lample and Conneau (2019) Guillaume Lample and Alexis Conneau. 2019. Cross-lingual Language Model Pretraining. CoRR abs/1901.07291 (2019). arXiv:1901.07291 http://arxiv.org/abs/1901.07291
@@
- Müller et al. (2020) Benjamin Müller, Antonis Anastasopoulos, Benoît Sagot, and Djamé Seddah. 2020. When Being Unseen from mBERT is just the Beginning: Handling New Languages With Multilingual Language Models. CoRR abs/2010.12858 (2020). arXiv:2010.12858 https://arxiv.org/abs/2010.12858
@@
- Nangia et al. (2020a) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel R. Bowman. 2020a. CrowS-Pairs: A Challenge Dataset for Measuring Social Biases in Masked Language Models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 1953–1967. https://doi.org/10.18653/v1/2020.emnlp-main.154
- Nangia et al. (2020b) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel R. Bowman. 2020b. CrowS-Pairs: A Challenge Dataset for Measuring Social Biases in Masked Language Models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 1953–1967. https://doi.org/10.18653/v1/2020.emnlp-main.154
@@
- OpenAI (2023) OpenAI. 2023. GPT-4 Technical Report. ArXiv abs/2303.08774 (2023). https://api.semanticscholar.org/CorpusID:257532815
@@
- Shi et al. (2022) Freda Shi, Mirac Suzgun, Markus Freitag, Xuezhi Wang, Suraj Srivats, Soroush Vosoughi, Hyung Won Chung, Yi Tay, Sebastian Ruder, Denny Zhou, Dipanjan Das, and Jason Wei. 2022. Language Models are Multilingual Chain-of-Thought Reasoners. arXiv:2210.03057 [cs.CL]
@@
- Thelwell (2020) Chinua Thelwell. 2020. Exporting Jim Crow: Blackface Minstrelsy in South Africa and Beyond. University of Massachusetts Press. http://www.jstor.org/stable/j.ctv160btb3
@@
- Weidinger et al. (2021) Laura Weidinger, John Mellor, Maribeth Rauh, Conor Griffin, Jonathan Uesato, Po-Sen Huang, Myra Cheng, Mia Glaese, Borja Balle, Atoosa Kasirzadeh, Zac Kenton, Sasha Brown, Will Hawkins, Tom Stepleton, Courtney Biles, Abeba Birhane, Julia Haas, Laura Rimell, Lisa Anne Hendricks, William Isaac, Sean Legassick, Geoffrey Irving, and Iason Gabriel. 2021. Ethical and social risks of harm from Language Models. arXiv:2112.04359 [cs.CL]
@@
- Xue et al. (2020) Linting Xue, Noah Constant, Adam Roberts, Mihir Kale, Rami Al-Rfou, Aditya Siddhant, Aditya Barua, and Colin Raffel. 2020. mT5: A massively multilingual pre-trained text-to-text transformer. CoRR abs/2010.11934 (2020). arXiv:2010.11934 https://arxiv.org/abs/2010.11934
@@ Appendix A Human Study
We followed the same approach as in (Cao et al., 2022) to collect human stereotypes. Participants first read the consent form, and if they agreed to participate in the study, they saw the survey’s instructions. For each social group, participants read in their respective language, “As viewed by American/Russian/Chinese/Indian society, (while my own opinions may differ), how [e.g., powerless, dominant, poor] versus [e.g., powerful, dominated, wealthy] are <group>?”.
They then rated each trait pair on a - slider scale representing the two poles of the trait pair (e.g. powerless and powerful). Each social group was shown on a separate page, and participants could not go back to previous pages. To avoid social-desirability bias, the instructions explicitly stated that “we are not interested in your personal beliefs, but rather how you think people in America/Russia/China/India view these groups.” Each participant was paid to rate social groups on pairs of traits and on average participants spent about minutes on the survey. This resulted in a pay of per hour. Maryland’s current minimum wage is 777https://www.minimum-wage.org/maryland. This study received the IRB approval.
A.1. Quality Assurance
Collecting high-quality data in subjective tasks is challenging since no ground truth exists. We followed the same quality control procedure as described in (Cao et al., 2022). Only crowd workers with an approval rate exceeding were eligible to participate in the survey. Each crowd worker had to successfully pass test questions in order for us to use their annotation888All participants were paid regardless of the quality check results..
For each group, we collected at least annotations that met our quality threshold. We collected annotations from a total of participants, out of which successfully passed the quality tests. We had participants that passed the quality tests for the English language, for Russian, for Chinese, and for Hindi. This indicated the significance of having such tests in place.
A.2. Participant Demographics
We collected participants’ demographic information including gender, age, education level, and (for non-English speakers) information about how frequently they read American social media. Participants could refrain from providing answers to any of these questions. After averaging the gender distribution across all languages: men , women , non-binary/transgender/gender fluid , and the rest of the participants preferred not to answer. Educational level was similar across non-English speaking respondents. On average, percent of respondents held a bachelor’s degree, master’s degree percent, Ph.D. , and the rest of the participants either preferred not to answer or held one of the following: associate degree, less than high-school graduate, professional degree (JD, MD, DVM, etc.). We didn’t have English-speaking respondents with a Ph.D., the percentage with a master’s degree was lower (), and the number of high-school graduates or equivalent was higher ().
For the English survey, the biggest ratio of annotators lived in Texas , for California and New York. The rest is distributed among states.
Age distribution for participants from all countries was more skewed towards younger people: on average, percent were between and years old, were between and years old, and the rest were older than . The youngest participant was years old and the oldest participant was years old.
Participants in the Russian survey were the ones who read American media most frequently: read it regularly compared to and percent for Hindi and Chinese respectively. On average, respondents read American media from time to time. Around never read the media.
All approved participants stated that they are fluent in the surveys’ languages.
Appendix B Related Works
The majority of studies on stereotypes in multilingual large language models (MLLMs) cover gender biases and use pairs of sentences translated into the subject languages (Cabello Piqueras and Søgaard, 2022; Wang et al., 2021; Kaneko et al., 2022; Steinborn et al., 2022; Bartl et al., 2020; Touileb et al., 2022). There are works, which use bias-prompting techniques and study how biases are expressed in different languages compared to English in domains of race, religion, ethnicity, and nationality (Levy et al., 2023; Câmara et al., 2022). According to Levy and colleagues (Levy et al., 2023), various languages result in distinct manifestations of biases. Camara and colleagues (Câmara et al., 2022) propose a framework to measure uni-sectional and intersectional biases across models trained on sentiment analysis tasks. There is work that compares how linguistically fair across different languages are multilingual models (Choudhury and Deshpande, 2021). Zhao and colleagues (Zhao et al., 2020) analyze bias in multilingual word embeddings and create a dataset in four languages. Numerous studies have put forth multilingual datasets for a wide range of tasks. Another work introduces a template-based anti-reflexive bias challenge dataset for Danish, Swedish, Chinese, and Russian languages that all have anti-reflexive gendered pronouns (González et al., 2020). Shi and colleagues developed a benchmark dataset for arithmetic reasoning in languages and showed that large pre-trained language models such as GPT3 are capable of performing multi-step reasoning across multiple languages (Shi et al., 2022). There is the CrowS dataset of sentence pairs in English for measuring bias in masked language models (Nangia et al., 2020b) and its extension to French language (Névéol et al., 2022).

### v3

@@ 4.2. Qualitative Results
We then delve into the specific stereotypical associations that transfer from one language to another, considering the potential impact of such strengthened associations. Our focus is on the GPT-3.5 model, where we observe the highest degree of stereotype leakage. For each source-target language pair that exhibits significant stereotype leakage based on our results, we examine the traits most strongly associated with each group in the target language. We pay particular attention to traits that are not deemed associated to the group according to target language human stereotypes but align with stereotypes from the source language. Our analysis identifies two primary types of leakage: the amplification of positive and negative associations for certain languages. Additionally, we observe non-polar leakages, which are characterized by associations that are neither positive nor negative.
@@ 4.2.1. Positive Leakage
According to human annotation, Asian people are more positively perceived in the English language than in Russian. We observe the strengthening of such traits in GPT-3.5 Russian language as wealthy, likable, and high status, possibly resulting from leakages from English and other languages. Moreover, housewives become more warm in English following leakages from possibly Russian and Hindi. Black people are more powerful, modern, confident, and wealthy in the English language following leakage from Hindi. Another example of the leakage of positive perceptions is for gay men and lesbians from English to other languages. Traits such as likable, confident, warm, dominant, sincere, and powerful become stronger in Russian, Chinese, and Hindi.
@@ 4.2.2. Negative Leakage
Meanwhile, there are negative stereotypes that leak across languages. From feminists, we observe a leakage from English to Chinese and Hindi, and from Russian to Chinese of such stereotypical associations as egoistic, threatening, repellent, and cold, while in the human data in Hindi, this group is perceived as warm.
Another example is immigrants. From Russian and English languages, traits such as threatening, repellent, dishonest, egoistic, and unconfident leak to Chinese and Hindi. Based on human data, we found that people surveyed in Chinese view this group quite favorably since the majority of immigrants to China were highly qualified professionals (Pieke, 2012). Contrarily, in Russia, immigrants are mostly coming from poorer neighboring countries and are negatively stereotyped in society, while in the U.S., immigrants are diverse and could be both marginalized or privileged.
@@ 4.2.3. Non-polar Leakage
There are also non-polar leakages, which are neither positive nor negative. From Hindi to English and Russian, we see the strengthening of religious for various groups such as women, disabled people, Black people, and Asian people. It has been shown that there are more than believers of the total population in India as of (Sahgal et al., 2021).
@@ 4.2.4. Non-shared Groups Leakage
In the case of non-shared groups, we expected uni-directional transferring of the groups’ perceptions from the language of origin to other languages. Our findings confirm this hypothesis. For example, the group VDV soldiers is a widely known military unit in Russia. There are strong stereotypes in Russian society about this group, but the group is mostly unknown to Americans. Out of the survey English survey respondents who passed the quality tests, no one chose this group as a familiar one. Stereotypes of this group leak from Russian to English, strengthening traits such as confident, traditional, competitive, and threatening. Another example is the Hui people, a group widely unknown to Russian and Hindi society: out of respondents for both surveys, no one chose this group as the familiar one. This social group is a minority in China and is composed of Chinese-speaking followers of Islam. Originally, Hui people were marginalized in China and viewed as more traditional, religious, and conservative (Hillman, 2004; Hong, 2005). Accordingly, we observed the leakage of such traits as irrational, traditional, threatening, repellent, religious, and egoistic. All groups specific to the Hindi language — Gujarati, Brahmin, and Shudra people — have certain traits leaking to the English and Russian languages. For example, high caste groups (Gujarati and Brahmin people) strengthen such positive traits as wealthy, likable, sincere, powerful, high status, competitive, and confident. In addition, Brahmin people become more associated in GPT-3.5 with traits poor, low status, powerless, traditional, religious, and dominated. This leakage corresponds to the perception of these groups in Indian society and by our survey respondents (Witzel, 1993; Milner, 1993).
4.3. Discussion
@@
However, positive stereotypes, such as the notion that Asian people are wealthy or housewives are warm, can also lead to unrealistic expectations and pressures. For instance, not all Asian people are wealthy, and assuming so can ignore the diverse economic realities faced by individuals within this broad demographic. Similarly, the stereotype that housewives are inherently warm can enforce restrictive roles based on gender.
@@
The leakage of stereotypes is particularly troubling for certain applications, including education and creative content generation. These fields heavily influence public perception and personal development, making the integrity of the content they deliver crucial. Systems built for these applications with MLLMs must be particularly cautious of the stereotype leakage effect. Developers need to implement strategies that actively mitigate the harmful leakage effects.
5. Conclusion
@@
Multilingual large language models have the potential to spread stereotypes beyond the societal context they emerge from, whether by generating new stereotypes, amplifying existing ones, or reinforcing prevailing social perceptions from dominant cultures. In our study, we demonstrate that this concern is indeed valid. To do so, we establish a framework for measuring the leakage of stereotypical associations in multilingual large language models across languages. Overall, we find that the stereotype leakage occurs bidirectionally meaning that when one language transmits stereotypes to others, it likely receives some stereotypes from other languages as well. We also observe the most stereotype leakage effect within the GPT-3.5 model.
Within the GPT-3.5 model, we observe the strengthening of positive, negative, and non-polar associations in the model. In addition, our study underscores the role of “native” languages in framing social groups unknown to other linguistic communities. Such leakage of stereotypes amplifies the complexity of societal perceptions by introducing a complex interconnected bias from different languages and cultures. In the context of shared groups, stereotype leakage may manifest as the manifestation of stereotypes that were not previously present within the cultural setting of a particular group. For non-shared groups, stereotype leakage can extend the reach of existing stereotypes from the source culture to other cultural contexts.
@@
To our knowledge, we are the first to introduce the concept of stereotype leakage across languages in multilingual LLMs. We propose a framework for quantifying this leakage in multilingual models, which can be easily applied to unstudied social groups. We show that multilingual large language models could facilitate the transmission of biases across different cultures and languages. We demonstrate the existence of stereotype leakage within MLLMs, which are trained on diverse linguistic datasets. As multilingual models begin to play an increasingly influential role in AI applications and across societies, understanding their potential vulnerabilities and the level of bias propagation across linguistic boundaries becomes important. As a result, we lay the groundwork for advancing both the theoretical comprehension of multilingual models and the practical implementation of bias mitigation in AI systems.
Limitations and Ethical Considerations
Our study has several limitations. First, we are limited in our ability to run a causal analysis because none of the studied languages can be easily removed from the training data to see their genuine impact on stereotypical associations in other languages. Retraining GPT-3.5, for instance, is not a feasible option. Thus, we use the BERT monolingual model as a proxy for each language.
In addition, stereotype traits were selected based on the ABC model, which was developed and tested using U.S. and German stereotypes. Though we translated our surveys into all four languages, the stereotype traits may better reflect Anglocentric stereotypes (Talat et al., 2022) than others.
Furthermore, the human stereotypes we collected may already reflect the influence of social stereotype transmission. For instance, in our study, we surveyed crowd workers about their consumption of U.S. social media. We found that, on average, of respondents from Russia, China, and India engage with U.S. social platforms. Such American cultural dominance could affect the human stereotypes collected in these three languages.
Lastly, while we indirectly consider culture through survey results on associations, we do not measure or account for culture comprehensively. Our English language survey results only apply to the U.S., Russian to Russia, Chinese to China, and Hindi to India.
@@ Acknowledgements.
- Abele et al. (2020) Andrea Abele, Naomi Ellemers, Susan Fiske, Alex Koch, and Vincent Yzerbyt. 2020. Navigating the social world: Toward an integrated framework for evaluating self, individuals, and groups. Psychological review 128 (09 2020). https://doi.org/10.1037/rev0000262
@@
- Dev et al. (2023) Sunipa Dev, Akshita Jha, Jaya Goyal, Dinesh Tewari, Shachi Dave, and Vinodkumar Prabhakaran. 2023. Building Stereotype Repositories with Complementary Approaches for Scale and Depth. In Proceedings of the First Workshop on Cross-Cultural Considerations in NLP (C3NLP), Sunipa Dev, Vinodkumar Prabhakaran, David Adelani, Dirk Hovy, and Luciana Benotti (Eds.). Association for Computational Linguistics, Dubrovnik, Croatia, 84–90. https://doi.org/10.18653/v1/2023.c3nlp-1.9
@@
- Koch et al. (2021) Alex Koch, Vincent Yzerbyt, Andrea Abele, Naomi Ellemers, and Susan T. Fiske. 2021. Social evaluation: Comparing models across interpersonal, intragroup, intergroup, several-group, and many-group contexts. Vol. 63. Elsevier, 1–68. https://doi.org/10.1016/bs.aesp.2020.11.001
@@
- Nangia et al. (2020) Nikita Nangia, Clara Vania, Rasika Bhalerao, and Samuel R. Bowman. 2020. CrowS-Pairs: A Challenge Dataset for Measuring Social Biases in Masked Language Models. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing (EMNLP). Association for Computational Linguistics, Online, 1953–1967. https://doi.org/10.18653/v1/2020.emnlp-main.154
@@
- Ouyang et al. (2022) Long Ouyang, Jeff Wu, Xu Jiang, Diogo Almeida, Carroll L. Wainwright, Pamela Mishkin, Chong Zhang, Sandhini Agarwal, Katarina Slama, Alex Ray, John Schulman, Jacob Hilton, Fraser Kelton, Luke Miller, Maddie Simens, Amanda Askell, Peter Welinder, Paul Christiano, Jan Leike, and Ryan Lowe. 2022. Training language models to follow instructions with human feedback. arXiv:2203.02155 [cs.CL]
@@ Appendix A Human Study
In this Section, we present details about the survey design, annotations quality control, and participants’ demographics.
A.1. Survey Design
Participants first reviewed a consent form, which outlined the purpose of the study, data usage, and confidentiality. Only after agreeing to participate, they proceed to the survey instructions. The consent form is shown in Figure 3
For each social group, participants read the following prompt in their respective language: “As viewed by American/Russian/Chinese/Indian society, (while my own opinions may differ), how [e.g., powerless, dominant, poor] versus [e.g., powerful, dominated, wealthy] are <group>?” They then rated each group on a slider scale ranging from -50 to 50, where the two poles of the scale represented opposite traits (e.g. powerless and powerful). Each social group appeared on a separate page, and participants were unable to return to previous pages, helping to minimize response bias. An example of the task is presented in Figure 4
To reduce social desirability bias, the instructions clearly emphasized: “We are not interested in your personal beliefs, but rather in how you think people in the United States/Russia/China/India view these groups.” The exact formulation is presented in Figure 5.
Participants were paid $2.00 to rate five social groups on 16 pairs of traits, which took an average of 10 minutes to complete, translating to a compensation rate of $12.00 per hour.
A.2. Quality Assurance
Collecting high-quality data for subjective tasks presents significant challenges, particularly due to the absence of objective ground truth. To mitigate these challenges, we implemented rigorous quality control procedures to ensure reliability and consistency across annotations.
The survey was administered through the Prolific platform, and only participants with an approval rate exceeding 90% were eligible to participate. This threshold was selected to balance data quality with participant availability, as it is generally considered high for Prolific, increasing the likelihood of obtaining reliable data.
In addition to the platform’s approval rate, we implemented three test questions throughout the survey to assess attentiveness and comprehension:
After the first group, participants must name the group they just scored.
After the second, participants must list one trait they just marked high and one marked low.
The fifth (final) group is a repetition of one of the four groups they previously scored.
We exclude annotators who answered the first two questions incorrectly. We then measured their intra-annotator (self) agreement by comparing the consistency of their responses, and any annotation with less than 80% self-agreement was discarded. These measures helped ensure data quality, though all participants were compensated regardless of their performance in the quality tests.
We collected at least five valid annotations per group that met our quality thresholds. Of the 286 participants, 151 passed the quality checks. Specifically, 34 participants passed for the English-language survey, 36 for Russian, 41 for Chinese, and 40 for Hindi. The fact that nearly half of the participants failed to meet the quality criteria underscores the necessity of these controls in subjective data collection.
A.3. Participant Demographics
We collected demographic information from participants, including gender, age, education level, and, for non-English speakers, their frequency of reading American social media. Participants were free to skip any question they preferred not to answer.
Across all languages, the gender distribution revealed a near balance: 49% identified as male, 45% as female, and 5% as non-binary, transgender, or gender fluid, with a few opting not to disclose. When we examined educational backgrounds, participants from non-English-speaking countries showed similar trends: 36% held a bachelor’s degree, 32% had a master’s degree, and 7% had earned a Ph.D. The remaining respondents either had lower educational qualifications or chose not to answer. English-speaking participants stood out, with no respondents holding a Ph.D., a lower percentage with master’s degrees (29%), and a larger proportion (35%) being high school graduates.
Among the English-speaking survey group, the largest proportion of respondents hailed from Texas (15%), followed by California and New York (each contributing 9%). The remaining participants were dispersed across 25 states, with no significant regional concentration outside these key areas.
As we looked at the age distribution, it was clear that younger people dominated the study, with 42% aged between 18 and 30 and 33% falling in the 31 to 40 range. The remainder were above 40, with the youngest participant being 18 and the oldest, a more experienced 72.
One notable demographic trend emerged in media consumption habits. Russian-speaking participants were the most frequent consumers of American media, with 44% stating they read it regularly. In contrast, 35% of Hindi-speaking participants and 28% of Chinese-speaking participants reported similar habits. Across all groups, about 39% said they occasionally consumed American media, while only 5% never did. These patterns suggest that Russian participants may be more exposed to or interested in global perspectives, particularly through American social media.
Crucially, all approved participants confirmed fluency in the language of their respective surveys. This ensures that any differences in responses were not influenced by language proficiency but more likely reflect deeper cultural or regional perspectives.
Appendix B Model Stereotypical Association Measurement
B.1. GPT Model Generation Failures on Marginalized Groups
We observe that for certain groups like feminist and Muslim person in Chinese, the model often disregards the prompt and simply outputs the group name. Moreover, in some cases, the model alters the trait specified in the prompt. For example, it changes dominating to dominated for disabled person in English or poor to wealthy for migrant worker in Russian. Additionally, the model may overlook the traits provided in the prompt and generate stereotypical traits instead. For instance, in Russian, it generates rape and patriot for Puerto Rican or cowboy for Texan.
We also count the number of system generations that did not match the instruction requirements for each social group. For example, in Chinese, we observed 108 generations for the group “feminist” that did not match the instruction requirement out of a total of 2880 generations. In comparison, there were 20 non-matching generations for “women.” However, these figures represent only an upper bound of system failures, as various reasons, such as generating synonyms, could cause mismatches. As stated in the paper, we leave the in-depth analysis for future work.